The gravitational force exerted on a solid object
Answer #1

F = mg

mass m = 5/9.8 = 0.51 kg

when submerged in water force of buoyant will act

and its magnitude is = (5 -3.5) = 1.5 N

Volume of water displaced is

density_water * V * g = F  

          V = 1.5 / 1000 * 9.8= 1.53 * 10-4 m3

This is equal to volume of object

Density of object = 0.51/ 1.53 * 10-4

                   = 3.33 * 103 kg/m3